OpenStudy (anonymous):

H5P1 ZERO-OFFSET AMPLIFIER In many amplifiers, we use dual power supplies so we can obtain a 0V offset at the output, that is, the DC operating point at the output is 0V. An example amplifier circuit that can achieve this is shown below. In this circuit, VS+=1.0V, VS−=−1.0V, and the MOSFET parameters are K=1 mAV2 and VT=0.5V. what formulas do i need

OpenStudy (anonymous):

tell me the formula

OpenStudy (nurali):

-0.5, 8000, 0.093

OpenStudy (nurali):

if correct please click best response

OpenStudy (anonymous):

H5P2 Source Follower Large Signal ....question 4

OpenStudy (anonymous):

i m not gettin the lab part of homework 5..........can u help me

OpenStudy (anonymous):

Run Transient Analysis. From plot, measure linear operating region: upper is 0.6V and lower is 1.16 V (approx). For Gain: Run Tran Analysis, measure output peak to peak= 1.94V (approx) and input peak to peak is 0.2V. So Gain = 1.94/0.2 = 9.7 For last part, there is a distortion. We know from Part 1 that the largest signal input that can be amplified without distortion is 1.16-0.6 = 0.56V, so amplitude is 0.56/2 = 0.28V